OpenStudy (anonymous):

Malcom has $1043 in bills of the following denominations: hundreds, fifties, twenties, fives, and ones. He has three times as many fifties as hundreds, twelve times as many twenties as hundreds, and four times as many ones as hundreds. He has nine more fives than hundreds. How many of each denomination of bill does he have?

OpenStudy (ranga):

Assume he has x hundred dollar bills. Then read the problem and see what they say about other bills. For example, "He has three times as many fifties as hundreds" means The number of 50 dollar bills will be 3x You try for the other bills.

OpenStudy (anonymous):

so it be 3x + 50 = $1054?

OpenStudy (anonymous):

what this mean? twelve times as many twenties as hundreds, and four times as many ones as hundreds. He has nine more fives than hundreds. 12x + 100 ~~~~~ 4x + 100s? 5x + 100?

OpenStudy (ranga):

"twelve times as many twenties as hundreds" means the number of 20 dollar bills that he has is 12 times the number of hundred dollar bills. We assumed he has x hundred dollar bills. So he must have 12x twenty dollar bills.

OpenStudy (anonymous):

OH i think i get it... 5x + 50 = 1054 12x + 100? = 1054

OpenStudy (ranga):

No. Each one of them does not add to $1043 No. of hundred dollar bills = x No. of fifty dollar bills = 3x No. of twenty dollar bills = 12x No. of one dollar bills = 4x x hundred dollar bills means, the amount is 100x 3x fifty dollar bills means, the amount is 50 times 3x = 150x 12x hundred dollar bills means, the amount is 20 times 12x = 240x 4x one dollar bills means, the amount is 1 times 4x = 4x Add up all the individual amounts and equate it to $1043 and solve for x

OpenStudy (anonymous):

ANSWER IS 1.94 ?

OpenStudy (anonymous):

u forgot 9 times 5x

OpenStudy (ranga):

Yes, No. of five dollar bills = (9 + x) For the five dollar bills it is different. It is not 9 times but 9 more. So it is (9+x). No. of hundred dollar bills = x No. of fifty dollar bills = 3x No. of twenty dollar bills = 12x No. of one dollar bills = 4x No. of five dollar bills = (9 + x) x hundred dollar bills means, the amount is 100x 3x fifty dollar bills means, the amount is 50 times 3x = 150x 12x hundred dollar bills means, the amount is 20 times 12x = 240x 4x one dollar bills means, the amount is 1 times 4x = 4x (9+x) five dollar bills means, the amount is 5 times (9+x) = 45 + 5x Add up all the individual amounts and equate it to $1043 and solve for x